At the rate 025 per unit of commodity x 250 unit commodity was bought given for certain reasons the price of commodity x increase in #50 per unit the quantity demanded of price was reduce to #150 compute the coefficient of cross price elasticity and state the relationship between the two commodity
1
Expert's answer
2019-05-27T11:57:19-0400

The coefficient of cross price elasticity is:

Ecp = (150 - 250)/(50 - 25)×(50 + 25)/(150 + 250) = -100/25×75/400 = -0.75.

So, these commodities are complements.